You have a deck of 52 playing cards.
(i) How many different 5 card hands can be dealt?
(ii) What is the probability that a hand of 5 dealt randomly contains 3 aces?
(iii) What is the probability that a hand of 5 dealt randomly will have 5 cards of the same
1
Expert's answer
2020-11-26T12:50:10-0500

i)To find the probability, we need to find the fraction where the numerator is the number of ways to have a flush and the denominator is the number of 5 card hands possible.

Each of these numbers will be found using Combinations

"C
n
,
k
=
\frac{n
!}{
(
k
)
!
(
n
−
k
)
!}"

"C
52
,
5
=
\frac{52
!}{
(
5
)
!
(
52
−
5
)
!}" "=\n2\n,\n598\n,\n960" ways


ii) To get exactly 3 aces, you need to choose 3 of the 4 aces and 2 of the other 48 cards. The number of ways to do that is

C(4,3)*C(48,2) = "\frac{4
!}{
(
3
)
!
(
4
−
3
)
!}"*"\frac{48
!}{
(
2
)
!
(
48
−
2
)
!}" =4512


Hence probability= "\frac{4512}{2598960}" = 0.00174



iii)Probability of getting a hand that has 5 cards of the same suit (flush, straight flush, royal flush.

Calculate all hands that involve a flush (so not just a flush but also a straight flush and royal flush) so that we'll be looking at any hand that has five cards of the same suit (with a suit having 13 cards in total). We can express getting this with:"C_{13,5}"

Keep in mind that there are 4 suits this can happen in, but we only want 1, and so we need to multiply by "C_{4,1}"

. Putting it together then, we get


"\frac{4!}{1!(4-1)!)}\frac{13!}{5!(13-5)!}" = 5148


The probability of getting a hand with a flush is:

= "\frac{5148}{
2598960}" ="00198\n ."